what is the midpoint between the points (1,2) and (5,-6) in the xy plane?

Answer by BigToosie(32)

The Midpoint formula (x1 + x2)/2 , (y1 + y2)/2
The midpoint between the points (1,2) and (5,-6)
x1 = 1
x2 = 5
y1 = 2
y2 = -6
(1 + 5)/2 , (2 - 6)/2

6/2 = 3 , -4/2 = -2
Therefore the midpoint is (3,-2)